INFRARED SPECTROSCOPY OF OH AND OH-C_{2}H_{2} EMBEDDED IN HELIUM NANODROPLETS

\underline{PAUL RASTON}, TAO LIANG, STEVEN D. FLYNN, ALEXANDER M. MORRISON, AND GARY E. DOUBERLY, Department of Chemistry, University of Georgia, Athens, GA 30602-2556.

The hydroxyl radical and its complex with acetylene have been characterized in superfluid helium nanodroplets by infrared depletion spectroscopy. For OH, two sharp Q(3/2) lines were observed, with a separation that is consistent with a fivefold increase in the parity (lambda type) doubling of the ^{2}Π_{3/2} state in helium droplets relative to the gas phase. This increase is rationalized in terms of the differences in the potential energy surfaces for He-OH(A′) and He-OH(A′′), which are most pronounced at around 90^{∘}. Switching to the T-shaped OH-C_{2}H_{2} complex, depletion signals corresponding to OH and CH stretching transitions were observed. The spectra reveal that the electronic angular momentum of OH in the complex is only partially quenched, and to a similar degree as observed in the gas phase^{a}^{,}^{b}. This indicates that the helium droplet environment does not significantly affect the electronic intermolecular interactions in OH-C_{2}H_{2}.

aJ. B. Davey, M. E. Greenslade, M. D. Marshall, M. I. Lester, and M. D. Wheeler J. Chem. Phys., 121, 3009 (2004).

bM. D. Marshall, J. B. Davey, M. E. Greenslade, and M. I. Lester J. Chem. Phys., 121, 5845 (2004).